# On Representation Complexity of Model-based and Model-free Reinforcement Learning

### v1

@@ 1 Introductions
In statistical machine learning, the efficiency of an algorithm is usually measured by sample complexity. For the comparison of model-based and model-free RL algorithms, many previous works also focus on their sample efficiency gap, and model-based algorithm usually enjoys a better sample complexity than model-free algorihms (Jin et al., 2018; Zanette and Brunskill, 2019; Tu and Recht, 2019; Sun et al., 2019). In general, the error of learning  [^1] For example, the performance difference between the learned policy and the optimal policy, which can be translated to sample complexity can be decomposed into three parts: optimization error, statistical error, and approximation error. Many previous efforts focus on optimization errors (Singh et al., 2000; Agarwal et al., 2020; Zhan et al., 2023) and statistical errors Kakade (2003); Strehl et al. (2006); Auer et al. (2008); Azar et al. (2017); Rashidinejad et al. (2022); Zhu and Zhang (2023), while approximation error has been less explored. Although a line of previous work studies the sample complexity additionally caused by the approximation error under the assumption of a bounded model misspecification error (Jin et al., 2019; Wang et al., 2020; Zhu et al., 2023b; Huang et al., 2021; Zhu et al., 2023a), they did not take a further or deeper step to study for what types of function classes (including transition kernel, reward function, $Q$-function, etc.), it is reasonable to assume a small model misspecification error.
In this paper, we study approximation errors through the lens of representation complexities. Intuitively, if a function (transition kernel, reward function, $Q$-function, etc.) has a low representation complexity, it would be relatively easy to learn it within a function class of low complexity and misspecification error, thus implying a better sample complexity. The previous work Dong et al. (2020b) studies a special class of MDPs with state space $\mathcal{S}=[0,1]$, action space $\mathcal{A}=\{0,1\}$, transition kernel piecewise linear with constant pieces and a simple reward function. By showing that the optimal $Q$-function requires an exponential number of linear pieces to approximate, they provide a concrete example that the $Q$-function has a much larger representation complexity than the transition kernel and reward function, which implies that model-based algorithms enjoy better sample complexity. However, the MDP class they study is restrictive. Thus, it is unclear whether it is a universal phenomenon that the underlying models have a lower representation complexity than other quantities such as $Q$-functions. Moreover, their metrics of measuring representation complexity, i.e., the number of pieces of piece-wise linear functions, is not fundamental, rigorous, or applicable to more general functions.
@@ 1.1 Related work
Many previous results imply that there exists the sample efficiency gap between model-based and model-free algorithms in various settings, including tabular MDPs (Strehl et al., 2006; Azar et al., 2017; Jin et al., 2018; Zanette and Brunskill, 2019), linear quadratic regulators (Dean et al., 2018; Tu and Recht, 2018; Dean et al., 2020), contextual decision processes with function approximation (Sun et al., 2019), etc. The previous work Dong et al. (2020b) compares model-based and model-free algorithms through the lens of the expressivity of neural networks. They claim that the expressivity is a different angle from sample efficiency. On the contrary, in our work, we posit the representation complexity as one of the important reasons causing the gap in sample complexity.
@@
The main result of our paper is that, in some cases, the underlying ground-truth dynamics (including the transition kernels and reward models) are easier to represent and thus learn, which inspires us to utilize the knowledge of the ground-truth model to boost the learning algorithms. This is consistent with the methods of many previous algorithms that exploit the dynamics to boost model-free quantities (Buckman et al., 2018; Feinberg et al., 2018; Luo et al., 2018; Janner et al., 2019), perform model-based planning (Oh et al., 2017; Weber et al., 2017; Chua et al., 2018; Wang and Ba, 2019; Piché et al., 2018; Nagabandi et al., 2018; Du and Narasimhan, 2019) or improve the learning procedure via various other approaches (Levine and Koltun, 2013; Heess et al., 2015; Rajeswaran et al., 2016; Silver et al., 2017; Clavera et al., 2018).
@@
A line of previous work (Jin et al., 2019; Wang et al., 2020; Zhu et al., 2023b) study the sample complexity of RL algorithms in the presence of model misspecification. This bridges the connection between the approximation error and sample efficiency. However, these works directly assume a small model misspecification error without further justifying whether it is reasonable. Our results imply that assuming a small error of transition kernel or reward function might be more reasonable than $Q$-functions. Many other works study the approximation error and expressivity of neural networks (Bao et al., 2014; Lu et al., 2017; Dong et al., 2020a; Lu et al., 2021). Instead, we study approximation error through circuit complexity, which provides a novel perspective and rigorous framework for future research.
@@
Circuit complexity is one of the most fundamental concepts in the theory of computer science (TCS) and has been long and extensively explored (Savage, 1972; Valiant, 1975; Trakhtenbrot, 1984; Furst et al., 1984; Hastad, 1986; Smolensky, 1987; Razborov, 1987, 1989; Boppana and Sipser, 1990; Arora and Barak, 2009). In this work, we first introduce circuit complexity to reinforcement learning to study the representation complexity of different functions including transition kernel, reward function and $Q$-functions, which bridges an important connection between TCS and RL.
@@ 1.2 Notations
Let $\lceil x\rceil$ denote the smallest integer greater than or equal to $x$, and let $\lfloor x\rfloor$ denote the greatest integer less than or equal to $x$. We use $\{0,1\}^{n}$ to denote the set of $n$-bits binary strings. Let $\delta_{x}$ denote the Dirac measure: $\delta_{x}(A)=\begin{cases}1,&~{}x\in A\\ 0,&~{}x\notin A\end{cases}$ and we use $\mathbbm{1}(\cdot)$ to denote the indicator function. Let $[n]$ denote the set $\{1,2,\dots,n\}$ and let =$\mathbb{N}=\{0,1,2,\ldots\}$ denote the set of all natural numbers.
@@ 2.3 Circuit complexity
Circuit complexity has been extensively explored in depth. In this section, we introduce concepts related to our results. One can refer to Arora and Barak (2009); Vollmer (1999) for more details.
Definition 1 (Boolean circuits, adapted from Definition 6.1, Arora and Barak (2009)).
@@
Definition 2 (Circuit computation).
@@
Definition 5 (Parity).
For every $n\in\mathbb{Z}_{+}$, the $n$-variable parity function $\mathsf{PARITY}_{n}:\{0,1\}^{n}\to\{0,1\}$ is defined by $\mathsf{PARITY}_{n}(x_{1},\dots,x_{n})=\sum_{i=1}^{n}x_{i}~{}\mod 2$. The parity function $\mathsf{PARITY}:\{0,1\}^{*}\to\{0,1\}$ is defined by
$\mathsf{PARITY}(w):=\mathsf{PARITY}_{|w|}(w),~{}\forall w\in\{0,1\}^{*}.$
@@ 3 Theoretical Results
The previous work Dong et al. (2020b) also conveyed similar messages. However, they only study a very special MDP while we study a more general class of MDPs. Moreover, Dong et al. (2020b) measures the representation complexity of functions by the number of pieces for a piecewise linear function, which is not rigorous and not applicable to more general functions.
To study the representation complexity of MDPs under a more rigorous and fundamental framework, we introduce circuit complexity to facilitate the study of representation complexity. In this work, we focus on circuits with constant-depth. We first show a warm-up example in Section 3.1, and then extend our results to a broader class of MDPs in Section 3.2.
@@ 3.1 Warm up example
Definition 6 (Parity MDP).
[…]
@@ 3.2 A broader family of MDPs
Definition 7 (Control function).
@@
[…]
@@
[…]
@@
[scale=0.8, every node/.style=transform shape, bit/.style=draw, rectangle, minimum size=1cm, text centered, node distance=0pt]
@@
[…]
[below=of ellipsis1, align=center, yshift=-0.7cm, xshift=0.2cm] Control Bits;
[below=of ellipsis2, align=center, yshift=-0.7cm] Representation Bits;
[decorate,decoration=brace,amplitude=10pt,mirror,raise=4pt,yshift=0pt] (bit1.south west) – (bitb.south east) node [black,midway,yshift=-0.8cm] ;
[decorate,decoration=brace,amplitude=10pt,mirror,raise=4pt,yshift=0pt] (bitb1.south west) – (bitbn.south east) node [black,midway,yshift=-0.8cm] ;
@@
Remark 1 (Control function and control bits).
@@
Remark 2 (Representation bits).
@@
Remark 3 (Condition function).
@@
Theorem 1 (Separation result of majority MDP, unconditioned setting).
@@
Theorem 2 (Separation result of majority MDP, conditioned setting).
@@
The proof of Theorem 1 and Theorem 2 are deferred to Section B.3 and Section B.2. In short, they imply that $\mathbb{T},r\in\mathbf{AC}_{0}$ and $Q^{*}\notin\mathbf{AC}_{0}$. In fact, we show that the value functions cannot be computed by a circuit with polynomial size and constant depth. Due to the relationship $Q^{*}(s,a)=r(s,a)+V^{*}(\mathbb{T}(s,a))$, we will treat these two functions synonymously and refer to them collectively as the "value function."
@@ 4 Experiments
Theorems 1 and 2 indicate that there exists a broad class of MDPs in which the transition functions and the reward functions have much lower circuit complexity than the optimal $Q$-functions (actually also value functions according to our proof for Majority MDPs). This observation, therefore, implies that value functions might be harder to approximate than transition functions and the reward functions, and gives rise to the following question:
@@
In this section, we seek to answer this question via experiments on common simulated environments. Specifically, we fix $d,m\in\mathbb{Z}_{+}$, and let $\mathcal{F}$ denote the class of $d$-depth, $m$-width neural networks (with input and output dimensions tailored to the context). The quantities of interest are the following relative approximation errors
$e_{\text{model}}=$ $~{}\min_{P\in\mathcal{F}}\frac{\mathbb{E}[\|P(s,a)-\mathbb{T}(s,a)\|^{2}]}{\mathbb{E}[\|\mathbb{T}(s,a)\|^{2}]}$
$e_{\text{reward}}=$ $~{}\min_{R\in\mathcal{F}}\frac{\mathbb{E}[(R(s,a)-r(s,a))^{2}]}{\mathbb{E}[(r(s,a))^{2}]}$
$e_{\text{$Q$-function}}=$ $~{}\min_{Q\in\mathcal{F}}\frac{\mathbb{E}[(Q(s,a)-Q^{*}(s,a))^{2}]}{\mathbb{E}[(Q^{*}(s,a))^{2}]}$
@@
Figure 3: Approximation errors of the optimal $Q$-functions, reward functions, and transition functions in MuJoCo environments. In each environment, we run $5$ independent experiments and report the mean and standard deviation of the approximation errors.
For common MuJoCo Gym environments (Brockman et al., 2016), including Ant-v4, Hopper-v4, HalfCheetah-v4, InvertedPendulum-v4, and Walker2d-v4 , we find these objectives by training $d$-depth, $m$-width neural networks to fit the corresponding values over the trajectories generated by an SAC-trained agent.  [^3] Our code is available at https://github.com/realgourmet/rep_complexity_rl. In Figure 3, we visualize [^4] We used Weights & Biases (Biewald, 2020) for experiment tracking and visualizations to develop insights for this paper. the approximation errors under $d=2,w=32$. Among the approximation objectives, the reward functions and the model transition functions are accessible, and we use Soft-Actor-Critic (Haarnoja et al., 2018) to learn the optimal $Q$-function. We observe a consistent phenomenon that the approximation errors of the optimal $Q$-function, in all environments, are much greater than the approximation errors of the transition and reward function. This finding concludes that in the above environments, the optimal $Q$-functions are more difficult to approximate than the transition functions and the reward function, which partly consolidate our hypothesis.
@@ Appendix A Supplementary Background on Circuit Complexity
$\mathsf{MAJORITY}(w):=\mathsf{MAJORITY}_{|w|}(w),~{}\forall w\in\{0,1\}^{*}.$
Proposition A.1 (Razborov [1987], Smolensky [1987, 1993]).
@@
$\mathsf{ADDITION}(w_{1},w_{2}):=\mathsf{ADDITION}_{\max\{|w_{1}|,|w_{2}|\}}(w_{1},w_{2}),~{}\forall w_{1},w_{2}\in\{0,1\}^{*}.$
Proposition A.2 (Proposition 1.15, Vollmer [1999]).
@@
$\mathsf{MAX}(w_{1},w_{2}):=\mathsf{MAX}_{\max\{|w_{1}|,|w_{2}|\}}(w_{1},w_{2}),~{}\forall w_{1},w_{2}\in\{0,1\}^{*}.$

### v2

@@ 1 Introductions
In statistical machine learning, the efficiency of an algorithm is usually measured by sample complexity. For the comparison of model-based and model-free RL algorithms, many previous works also focus on their sample efficiency gap, and model-based algorithm usually enjoys a better sample complexity than model-free algorihms (Jin et al., 2018; Zanette & Brunskill, 2019; Tu & Recht, 2019; Sun et al., 2019). In general, the error of learning  [^1] For example, the performance difference between the learned policy and the optimal policy, which can be translated to sample complexity. can be decomposed into three parts: optimization error, statistical error, and approximation error. Many previous efforts focus on optimization errors (Singh et al., 2000; Agarwal et al., 2020; Zhan et al., 2023) and statistical errors Kakade (2003); Strehl et al. (2006); Auer et al. (2008); Azar et al. (2017); Rashidinejad et al. (2022); Zhu & Zhang (2024), while approximation error has been less explored. Although a line of previous work studies the sample complexity additionally caused by the approximation error under the assumption of a bounded model misspecification error (Jin et al., 2019; Wang et al., 2020; Zhu et al., 2023; Huang et al., 2021; Zhu et al., 2024), they did not take a further or deeper step to study for what types of function classes (including transition kernel, reward function, $Q$-function, etc.), it is reasonable to assume a small model misspecification error.
In this paper, we study approximation errors through the lens of representation complexities. Intuitively, if a function (transition kernel, reward function, $Q$-function, etc.) has a low representation complexity, it would be relatively easy to learn it within a function class of low complexity and misspecification error, thus implying a better sample complexity. The previous work Dong et al. (2020) studies a special class of MDPs with state space $\mathcal{S}=[0,1]$, action space $\mathcal{A}=\{0,1\}$, transition kernel piecewise linear with constant pieces and a simple reward function. By showing that the optimal $Q$-function requires an exponential number of linear pieces to approximate, they provide a concrete example that the $Q$-function has a much larger representation complexity than the transition kernel and reward function, which implies that model-based algorithms enjoy better sample complexity. However, the MDP class they study is restrictive. Thus, it is unclear whether it is a universal phenomenon that the underlying models have a lower representation complexity than other quantities such as $Q$-functions. Moreover, their metrics of measuring representation complexity, i.e., the number of pieces of piece-wise linear functions, is not fundamental, rigorous, or applicable to more general functions.
@@ 1.1 Related work
Many previous results imply that there exists the sample efficiency gap between model-based and model-free algorithms in various settings, including tabular MDPs (Strehl et al., 2006; Azar et al., 2017; Jin et al., 2018; Zanette & Brunskill, 2019), linear quadratic regulators (Dean et al., 2018; Tu & Recht, 2018; Dean et al., 2020), contextual decision processes with function approximation (Sun et al., 2019), etc. The previous work Dong et al. (2020) compares model-based and model-free algorithms through the lens of the expressivity of neural networks. They claim that the expressivity is a different angle from sample efficiency. On the contrary, in our work, we posit the representation complexity as one of the important reasons causing the gap in sample complexity.
@@
The main result of our paper is that, in some cases, the underlying ground-truth dynamics (including the transition kernels and reward models) are easier to represent and thus learn, which inspires us to utilize the knowledge of the ground-truth model to boost the learning algorithms. This is consistent with the methods of many previous algorithms that exploit the dynamics to boost model-free quantities (Buckman et al., 2018; Feinberg et al., 2018; Luo et al., 2018; Janner et al., 2019), perform model-based planning (Oh et al., 2017; Weber et al., 2017; Chua et al., 2018; Wang & Ba, 2019; Piché et al., 2018; Nagabandi et al., 2018; Du & Narasimhan, 2019) or improve the learning procedure via various other approaches (Levine & Koltun, 2013; Heess et al., 2015; Rajeswaran et al., 2016; Silver et al., 2017; Clavera et al., 2018).
@@
A line of previous work (Jin et al., 2019; Wang et al., 2020; Zhu et al., 2023) study the sample complexity of RL algorithms in the presence of model misspecification. This bridges the connection between the approximation error and sample efficiency. However, these works directly assume a small model misspecification error without further justifying whether it is reasonable. Our results imply that assuming a small error of transition kernel or reward function might be more reasonable than $Q$-functions. Many other works study the approximation error and expressivity of neural networks (Bao et al., 2014; Lu et al., 2017; Dong et al., 2020; Lu et al., 2021). Instead, we study approximation error through circuit complexity, which provides a novel perspective and rigorous framework for future research.
@@
Circuit complexity is one of the most fundamental concepts in the theory of computer science (TCS) and has been long and extensively explored (Savage, 1972; Valiant, 1975; Trakhtenbrot, 1984; Furst et al., 1984; Hastad, 1986; Smolensky, 1987; Razborov, 1987; 1989; Boppana & Sipser, 1990; Arora & Barak, 2009). In this work, we first introduce circuit complexity to reinforcement learning to study the representation complexity of different functions including transition kernel, reward function and $Q$-functions, which bridges an important connection between TCS and RL.
@@ 1.2 Notations
Let $\lceil x\rceil$ denote the smallest integer greater than or equal to $x$, and let $\lfloor x\rfloor$ denote the greatest integer less than or equal to $x$. We use $\{0,1\}^{n}$ to denote the set of $n$-bits binary strings. Let $\delta_{x}$ denote the Dirac measure: $\delta_{x}(A)=\begin{cases}1,&\leavevmode\nobreak\ x\in A\\ 0,&\leavevmode\nobreak\ x\notin A\end{cases}$ and we use $\mathbbm{1}(\cdot)$ to denote the indicator function. Let $[n]$ denote the set $\{1,2,\dots,n\}$ and let =$\mathbb{N}=\{0,1,2,\ldots\}$ denote the set of all natural numbers.
@@ 2.3 Circuit complexity
Circuit complexity has been extensively explored in depth. In this section, we introduce concepts related to our results. One can refer to Arora & Barak (2009); Vollmer (1999) for more details.
Definition 2.1 (Boolean circuits, adapted from Definition 6.1, Arora & Barak (2009)).
@@
Definition 2.2 (Circuit computation).
@@
Definition 2.5 (Parity).
For every $n\in\mathbb{Z}_{+}$, the $n$-variable parity function $\mathsf{PARITY}_{n}:\{0,1\}^{n}\to\{0,1\}$ is defined by $\mathsf{PARITY}_{n}(x_{1},\dots,x_{n})=\sum_{i=1}^{n}x_{i}\leavevmode\nobreak\ \mod 2$. The parity function $\mathsf{PARITY}:\{0,1\}^{*}\to\{0,1\}$ is defined by
$\mathsf{PARITY}(w):=\mathsf{PARITY}_{|w|}(w),\leavevmode\nobreak\ \forall w\in\{0,1\}^{*}.$
@@ 3 Theoretical Results
The previous work Dong et al. (2020) also conveyed similar messages. However, they only study a very special MDP while we study a more general class of MDPs. Moreover, Dong et al. (2020) measures the representation complexity of functions by the number of pieces for a piecewise linear function, which is not rigorous and not applicable to more general functions.
To study the representation complexity of MDPs under a more rigorous and fundamental framework, we introduce circuit complexity to facilitate the study of representation complexity. In this work, we focus on circuits with constant-depth. We first show a warm-up example in Section 3.1, and then extend our results to a broader class of MDPs in Section 3.2. In Appendix E, we also briefly discuss possible extensions to POMDP and stochastic MDP.
@@ 3.1 Warm up example
Definition 3.1 (Parity MDP).
An $n$-bits parity MDP is defined as follows: the state space is $\{0,1\}^{n}$, the action space is $\{(i,j):i,j\in[n]\}$, and the planning horizon is $H=n$. Let the reward function be defined by: $r(s,a)=\begin{cases}1,&\leavevmode\nobreak\ s=\mathbf{0}_{n}\\ 0,&\leavevmode\nobreak\ \text{otherwise}\end{cases}$. Let the transition function be defined as follows: for each state $s$ and action $a=(i,j)$, transit with probability $1$ to $s^{\prime}$ where $s^{\prime}$ is given by flipping the $i$-th and $j$-th bits of $s$.
@@ 3.2 A broader family of MDPs
Definition 3.2 (Control function).
@@
  • The reward function is defined by: $r(s,a)=\begin{cases}1,&\leavevmode\nobreak\ s[r]=s_{\mathrm{reward}},s[c]=\mathbf{1}_{b}\\ 0,&\leavevmode\nobreak\ \text{otherwise}\end{cases}$;
@@
    $\mathbb{T}(s,a)=\begin{cases}\left(s[c],g(s[r],\sum_{j=1}^{b}2^{b-j}\cdot s_{j})\right),&\leavevmode\nobreak\ a=1,C(s[r])=1\\ \left(f(s[c]),s[r]\right),&\leavevmode\nobreak\ \text{otherwise}\end{cases}$
@@
Remark 3.4 (Control function and control bits).
@@
Remark 3.5 (Representation bits).
@@
Remark 3.6 (Condition function).
@@
Theorem 3.7 (Separation result of majority MDP, unconditioned setting).
@@
Theorem 3.8 (Separation result of majority MDP, conditioned setting).
@@
The proof of Theorem 3.7 and Theorem 3.8 are deferred to Section B.3 and Section B.2. In short, they imply that $\mathbb{T},r\in\mathbf{AC}_{0}$ and $Q^{*}\notin\mathbf{AC}_{0}$. In fact, we show that the value functions cannot be computed by a circuit with polynomial size and constant depth. Due to the relationship $Q^{*}(s,a)=r(s,a)+V^{*}(\mathbb{T}(s,a))$, we will treat these two functions synonymously and refer to them collectively as the "value function."
@@ 4 Experiments
Theorems 3.7 and 3.8 indicate that there exists a broad class of MDPs in which the transition functions and the reward functions have much lower circuit complexity than the optimal $Q$-functions (actually also value functions according to our proof for Majority MDPs). This observation, therefore, implies that value functions might be harder to approximate than transition functions and the reward functions, and gives rise to the following question:
@@
In this section, we seek to answer this question via experiments on common simulated environments. Specifically, we fix $d,m\in\mathbb{Z}_{+}$, and let $\mathcal{F}$ denote the class of $d$-depth (i.e., $d$ hidden layers), $m$-width neural networks (with input and output dimensions tailored to the context). The quantities of interest are the following relative approximation errors
$e_{\text{model}}=$ $\leavevmode\nobreak\ \min_{P\in\mathcal{F}}\frac{\mathbb{E}[\|P(s,a)-\mathbb{T}(s,a)\|^{2}]}{\mathbb{E}[\|\mathbb{T}(s,a)\|^{2}]}$
$e_{\text{reward}}=$ $\leavevmode\nobreak\ \min_{R\in\mathcal{F}}\frac{\mathbb{E}[(R(s,a)-r(s,a))^{2}]}{\mathbb{E}[(r(s,a))^{2}]}$
$e_{\text{$Q$-function}}=$ $\leavevmode\nobreak\ \min_{Q\in\mathcal{F}}\frac{\mathbb{E}[(Q(s,a)-Q^{*}(s,a))^{2}]}{\mathbb{E}[(Q^{*}(s,a))^{2}]}$
@@
Figure 3: Approximation errors of the optimal $Q$-functions, reward functions, and transition functions in MuJoCo environments. In each environment, we run $5$ independent experiments and report the mean and standard deviation of the approximation errors. All curves (as well as those in Figure 4-9) are displayed with an exponential average smoothing with rate $0.2$.
For common MuJoCo Gym environments (Brockman et al., 2016), including Ant-v4, Hopper-v4, HalfCheetah-v4, InvertedPendulum-v4, and Walker2d-v4 , we find these objectives by training $d$-depth, $m$-width neural networks to fit the corresponding values over the trajectories generated by an SAC-trained agent.  [^3] Our code is available at https://github.com/realgourmet/rep_complexity_rl. In Figure 3, we visualize [^4] We used Weights & Biases (Biewald, 2020) for experiment tracking and visualizations. the approximation errors under $d=2,w=32$. Among the approximation objectives, the reward functions and the model transition functions are accessible, and we use Soft-Actor-Critic (Haarnoja et al., 2018) to learn the optimal $Q$-function. We consistently observe that the approximation errors of the optimal $Q$-function, in all environments, are much greater than the approximation errors of the transition and reward function. This finding concludes that in the above environments, the optimal $Q$-functions are more difficult to approximate than the transition functions and the reward function, which partly consolidate our hypothesis. We provide experimental details in Appendix C and additional experiments in Appendix D.
@@ Appendix A Supplementary Background on Circuit Complexity
$\mathsf{MAJORITY}(w):=\mathsf{MAJORITY}_{|w|}(w),\leavevmode\nobreak\ \forall w\in\{0,1\}^{*}.$
Proposition A.2 (Razborov (1987); Smolensky (1987; 1993)).
@@
$\mathsf{ADDITION}(w_{1},w_{2}):=\mathsf{ADDITION}_{\max\{|w_{1}|,|w_{2}|\}}(w_{1},w_{2}),\leavevmode\nobreak\ \forall w_{1},w_{2}\in\{0,1\}^{*}.$
Proposition A.4 (Proposition 1.15, Vollmer (1999)).
@@
$\mathsf{MAX}(w_{1},w_{2}):=\mathsf{MAX}_{\max\{|w_{1}|,|w_{2}|\}}(w_{1},w_{2}),\leavevmode\nobreak\ \forall w_{1},w_{2}\in\{0,1\}^{*}.$